OpenStudy (anonymous):

Tutorial: How you can verify a scientist's claim in 10 minutes with the help of a scientific calculator.

OpenStudy (anonymous):

------ Anecdote: I was reading this book called "The Feynman Lectures on physics". As he talks about the electrical force, he goes "And all matter is a mixture of positive protons and negative electrons which are attracting and repelling with this great force. So perfect is the balance, however, that when you stand near someone else you don't feel any force at all. If there were even a little bit of unbalance, you would know it. If you were standing at arm's length from someone and each of you had one percent more electrons than protons, the repelling force would be incredible. How great? The repulsion would be enough to lift a "weight" equal to that of the entire earth!" And I go "Orly?" ------

OpenStudy (anonymous):

The number crunching: Let's say I weigh 70 kg* (154 lb). Let's assume that the human body is 70% water & 25% carbon. \(\Rightarrow \begin{array}{ll}\text{The mass of water}=70*0.7=49 kg. \\ \text{The mass of carbon}\approx 21 kg. \end{array} \\ \therefore\begin{array}{ll} \text{The no. of moles of water = 49/0.018 = 2722.2 } \\ \text{The no. of moles of carbon = 21/0.012=1750}\end{array}\) Remember that each mole of a substance contains \(6.022*10^{23}\) molecules. And each molecule of water & carbon has 18 & 12 electrons/protons respectively. (Considering a neutron to be made up of an electron & a proton). Therefore, the no. of electrons in the water=\(2722.2 * 6.022*10^{23} * 18=2.95*10^{28}\) And that in the carbon=\(1750 * 6.022*10^{23} * 12=1.26*10^{28}\) So, the total no. of electrons in my body=\((2.95*10^{28})+(1.26*10^{28})=4.21*10^{28}\) The excess electrons needed in each body = 1% of the above = \(4.21*10^{26}\) The excess charge (Q) in each body is then = n*e =\(4.21*10^{26} * 1.6*10^{-19}=6.736*10^{7}C\) If you are still with me, let's at last calculate the repulsive force between me an my imaginary twin brother (who is also charged to the same extent). I'll assume an "arm's length" to be 2.5 feet or 0.762m. Let us plug this data into the equation: \[\Large\begin{align}F&=\frac{(9*10^{9}) Q^{2}}{R^{2}}\\ F&=\frac{(9*10^{9})(6.736*10^{7})^{2}}{0.762^{2}}\\ F&\approx7*10^{25}N.\end{align}\] The mass of the earth is \(6*10^{24}\) kg. He did say that the force would be equal to that of the weight of the earth sooo... \(\Large\text{The weight of the earth is }\approx 6*10^{25}N.\) That's close enough for me. And so....*drumroll*....\[\color{DarkGreen} {\Large\text{HE WAS RIGHT!!}}\]

OpenStudy (anonymous):

*A misnomer. My mass is 70 kg.

OpenStudy (anonymous):

yepp, electrical forces are way too higher than gravitational forces

OpenStudy (anonymous):

Feynman had a nasty habit of being right a lot ;)

OpenStudy (anonymous):

By the way - he solved Engineering problems of equal difficulty to the scientific ones.

OpenStudy (shivaniits):

but the atomic mass of water and carbon is 18 and 12 and if i am considering neutron=electron+proton then 18=no of protons+no of neutrons further 18=no of protons+no of electrons+no of protons then 18=2x+y taking x as no of protons and y as no of electrons and now there are equal no of protons and electrons in atom then x=y putting it in eqn i have 3x=18 giving x=y=6 so no of electrons are 6 then how are you considering them to be 18....??

OpenStudy (anonymous):

An atom of Carbon has 6 electrons, 6 protons & 6 neutrons. If I consider the neutron to be made up of a proton & an electron, then I can consider an atom of Carbon to have a total of 12 electrons (& 12 protons & no neutrons).

OpenStudy (shivaniits):

got it!!